# Generalized Moving Least-Squares for Solving Vector-valued PDEs on Unknown Manifolds

### v1

@@ Abstract
In this paper, we extend the Generalized Moving Least-Squares (GMLS) method in two different ways to solve the vector-valued PDEs on unknown smooth 2D manifolds without boundaries embedded in $\mathbb{R}^{3}$, identified with randomly sampled point cloud data. The two approaches are referred to as the intrinsic method and the extrinsic method. For the intrinsic method which relies on local approximations of metric tensors, we simplify the formula of Laplacians and covariant derivatives acting on vector fields at the base point by calculating them in a local Monge coordinate system. On the other hand, the extrinsic method formulates tangential derivatives on a submanifold as the projection of the directional derivative in the ambient Euclidean space onto the tangent space of the submanifold. One challenge of this method is that the discretization of vector Laplacians yields a matrix whose size relies on the ambient dimension. To overcome this issue, we reduce the dimension of vector Laplacian matrices by employing an appropriate projection. The complexity of both methods scales well with the dimension of manifolds rather than the ambient dimension. We also present supporting numerical examples, including eigenvalue problems, linear Poisson equations, and nonlinear Burgers’ equations, to examine the numerical accuracy of proposed methods on various smooth manifolds.
@@ 1 Introduction
Solving vector-valued Partial Differential Equations (PDEs) on curved surfaces or embedded manifolds has various applications in natural sciences and practical engineering. Examples include transport processes of surfactants within thin films [23], surface dynamics of fluid vesicles and biomembranes [48, 4], formation of fractures and barriers on interfaces [35], tracking evolving geometries or even topological changes of interfaces [47], fluid flows in computer graphics [51], etc.
@@
For such wide applications, there have been several developing numerical methods to approximate the solution of vector-valued PDEs on manifolds. Given a sufficient high-quality simplicial complex or mesh, both Discrete Exterior Calculus (DEC) [24, 10] and Finite Element Exterior Calculus (FEEC) [2, 25, 18] can achieve the desirable convergence in PDE problems for manifolds embedded in a low-dimensional ambient space. However, constructing meshes on surfaces with high quality can be difficult and complicated when the manifold is only identified by randomly sampled data.
@@
To avoid this issue, various mesh-free approaches have been developed for solving PDEs in recent decades. For unknown manifolds with only a point cloud given, the first step of most mesh-free approaches is to characterize the manifolds using a parametrization scheme, such as local SVD methods [12, 60, 57], level set methods [6, 19, 59, 1], closest point methods [45, 41], orthogonal gradient methods [42], moving least squares [33], just to name a few. The second step is to approximate the differential operators acting on functions in the strong form along the tangent bundles using approaches including the Radial Basis Function (RBF) methods [15, 16], Generalized Moving Least-Squares (GMLS) [33, 20], Generalized Finite Difference Methods (GFDM) [54, 52]. In our previous work [22], the RBF method was extended to provide an accurate approximation of differential operators acting on vector fields of smooth manifolds. However, its effectiveness depends crucially on the choice of shape parameters in the radial basis function. Additionally, the computational cost of the method is high since the resulting discrete operators are dense matrices with their size dependent on the ambient dimension.
@@
Alternatively, graph-based approaches, including Vector Diffusion Maps (VDM) [50] and Spectral Exterior Calculus (SEC) [5] can also be used as a mesh-free tool to approximate vector Laplacians for processing raw data in manifold learning applications. This type of approaches does not need a parametrization of a point cloud while restricted to approximate a certain type of differential operators.
@@
In this paper, we consider solving vector-valued PDEs on two-dimensional unknown manifolds $M$ without boundaries embedded in the Euclidean space $\mathbb{R}^{3}$, identified with randomly sampled point cloud data. Our goal here is to construct an approximation for differential operators acting on vector fields of smooth manifolds where the resulting matrices are sparse with size only dependent of the intrinsic dimension. Based on two different types of formulations of vector differential operators, we extend the GMLS methods in two different ways to approximate the vector differential operators and subsequently solve vector-valued PDEs on manifolds. The two approximations will be referred to as the intrinsic method and the extrinsic method.
@@
For the intrinsic GMLS method, it in general involves with the local characterization of the Riemannian metric tensor and other geometric quantities of manifolds in local intrinsic coordinates [33, 20, 27]. In this paper, we derive the formula of the covariant derivative and Laplacians acting on vector fields at the base point in a local normal coordinate system. It turns out that the resulting formula can be much simplified by using the properties of the normal coordinate system so that it helps to reduce the computational cost in constructing the vector Laplacian matrices. Our derivations can be viewed as a generalization from the Laplace-Beltrami of functions in [33] to the covariant derivative and Laplacians of vector fields.
@@
For the extrinsic GMLS method, the surface derivative in the local coordinate can be written as a projection of the appropriate derivative in the ambient space [53, 52, 26]. Then the extrinsic method relies on the discretization in the ambient space which can be computationally expensive for the vector case. In this paper, we apply an appropriate projection to reduce the size of vector Laplacian matrices from $3N\times 3N$ to $2N\times 2N$, where $N$ is the point cloud data size, $3$ corresponds to the ambient dimension and $2$ corresponds to the intrinsic dimension of manifolds. This dimension reduction can be numerically realized since all objects (e.g., functions and tensor fields) in this paper are confined to the tangent bundle of manifolds and the surfaces are all static without movement in the normal direction. In fact, this work improves the result in our previous work [22], saving the memory cost and computational cost especially when the ambient dimension $n\sim 10-100$ is moderately large.
@@
The paper is organized as follows. In Section 2, we provide a short review of the GMLS approach and its application in the approximation of tangent vectors on unknown manifolds identified by randomly sampled data points. In Section 3, we present the intrinsic formulation for approximating the Bochner Laplacian and the covariant derivative using GMLS. In Section 4, using the extrinsic formulation, we apply a projection technique to improve the approximation for the Bochner Laplacian and the covariant derivative. In Section 5, we provide supporting numerical examples for solving vector-valued PDEs, particularly eigenvalue problems, Poisson equations, linear vector diffusion equations and nonlinear Burgers’ equations. In Section 6, we conclude the paper with a summary and some open problems. We relegate the intrinsic approximation and extrinsic approximation of two other vector Laplacians to the Appendix A and Appendix B, respectively.
@@ 2.1 Review of GMLS
The GMLS approach is a regression technique for approximating functions by a linear combination of local basis functions, such as local polynomials, based on a set of scattered data samples (see e.g. [40, 34, 32, 58, 37]). We now briefly review the application of GMLS approach on a 2D smooth manifold embedded in $\mathbb{R}^{3}$. Given a set of (distinct) nodes $\mathbf{X}_{M}=\{\mathbf{x}_{i}\}_{i=1}^{N}\subset M$. For a base point ${\mathbf{x}_{0}}\in\mathbf{X}_{M}\subset M$, we denote its $K$-nearest neighbors in $\mathbf{X}_{M}$ by $S_{{\mathbf{x}_{0}}}=\{{\mathbf{x}_{0,k}}\}_{k=1}^{K}\subset\mathbf{X}_{M}$. In most literature, such a set $S_{{\mathbf{x}_{0}}}$ is called a “stencil”. By definition, we have ${\mathbf{x}_{0,1}}={\mathbf{x}_{0}}$ to be the base point. We assume that the true tangent space of the manifold $M$ is given at this moment. We will discuss the unknown manifold regime in the following subsection 2.2. Let $T_{\mathbf{x}}M$ be the tangent space at point $\mathbf{x}=(x^{1},x^{2},x^{3})\in M$ and denote its orthonormal basis by $\left\{\bm{t}_{1}(\mathbf{x}),\bm{t}_{2}(\mathbf{x})\right\}$.
Figure 1: Left panel: Sketch of our setups. (a) Global tangent-vector bases $\{\bm{t}_{1},\bm{t}_{2}\}$ on the point cloud $\{\mathbf{x}_{i}\}_{i=1}^{N}$ (gray dots) without alignment. (b) $k$-nearest neighbors of the base point $\mathbf{x}_{0}$ plotted as blue crosses inside the red circle. (c) The tangent space $\mathrm{span}\{\bm{t}_{1},\bm{t}_{2}\}$ and the normal $\bm{n}$ at the base point $\mathbf{x}_{0}$. Also plotted are their estimations $\{\bm{\hat{t}}_{1},\bm{\hat{t}}_{2},\bm{\hat{n}}\}$ given in (8) when the manifold is unknown. (d) The local Monge coordinate $(\theta_{1},\theta_{2})$ and the GMLS approximation of the manifold $\hat{q}(\theta_{1},\theta_{2})$ in the local coordinate chart (equation (9)). Right panel: Convergence results for the estimations of normal vectors on a torus, identified by randomly sampled data points.
@@
where $\lambda_{k}$ is some positive weight function. There are various choices of weighted $\ell^{2}$-norms in the moving least-squares (2) as shown in [58, 33, 20]. As reported in previous works [33, 26], the weight function
@@
can numerically provide a stable approximation to the Laplace-Beltrami operator in many cases for randomly sampled nodes. Here we also apply such a special weight function for the stable vector Laplacians. For the covariant derivative, we choose the uniform weight function for a stable approximation.
@@
We now present the theoretical error estimates [37, 36] for such a GMLS approximation for partial derivative operators. For a set of points $\mathbf{X}_{M}=\{\mathbf{x}_{i}\}_{i=1}^{N}\subset M$, define the fill distance $h_{\mathbf{X},M}$ and the separation distance $q_{\mathbf{X},M}$ in Euclidean space as
@@
respectively, and assume that $q_{\mathbf{X},M}\leq h_{\mathbf{X},M}\leq c_{q}q_{\mathbf{X},M}$ for some constant $c_{q}>0$. Let $\mathbf{X}_{M}\subseteq M$ be such that $h_{\mathbf{X},M}\leq h_{0}$ for some constant $h_{0}>0$ and $M^{\ast}=\bigcup_{\mathbf{x}\in M}B(\mathbf{x},C_{2}h_{0})$, a union of geodesic ball over the length $C_{2}h_{0}$ with some constant $C_{2}>0.$ Then for any $f\in C^{l+1}(M^{\ast})$ and $\bm{\delta}=(\delta_{1},\delta_{2})$ that satisfies $|{\bm{\delta}}|=\delta_{1}+\delta_{2}\leq l$,
@@
for all $\mathbf{x}\in M$ and some $C>0$ [37]. In the error bound above, we used the notation $\widehat{D^{{\bm{\delta}}}f}$ as the GMLS approximation to $D^{{\bm{\delta}}}f$ using local polynomials up to degree $l$, where $D^{{\bm{\delta}}}$ denotes a general multi-dimensional derivative with multi-index ${\bm{\delta}}$. For uniformly i.i.d. random data $\mathbf{X}_{M}\subseteq M$ on manifolds, with probability higher than $1-\frac{1}{N}$, one can show that $h_{\mathbf{X},M}=O((\log N/N)^{1/d})$, where $d$ is the intrinsic dimension [26]. In the case of a 2D manifold, we have that $h_{\mathbf{X},M}=O(N^{-1/2})$ and consequently
$|D^{{\bm{\delta}}}f(\mathbf{x})-\widehat{D^{{\bm{\delta}}}f}(\mathbf{x})|=O(N^{-\frac{l+1-|{\bm{\delta}}|}{2}}),$ (7)
where we have ignored the $\log$ factor.
@@ 2.2 Approximation of tangent vectors for unknown manifolds
For the unknown manifold setup, we assume that we are only given a point cloud $\{\mathbf{x}_{i}\}_{i=1}^{N}\subset M$ on the interior of the smooth manifold. To construct approximations of differential operators, we need to approximate bases of tangent vector space $\left\{\bm{t}_{1}(\mathbf{x}),\bm{t}_{2}(\mathbf{x})\right\}$ on each point $\{\mathbf{x}_{i}\}_{i=1}^{N}$. We first use the local singular value decomposition (SVD) or local principal component analysis (PCA) for a coarse approximation of the tangent vectors [12, 60, 57, 33, 22]:
@@
    Take a singular value decomposition of $\mathbf{D}=\mathbf{V}_{L}\bm{\Sigma}\mathbf{V}_{R}$, and then obtain the leading 2-columns of $\mathbf{V}_{L}\in\mathbb{R}^{3\times 3}$, denoted by $\left\{\bm{\tilde{t}}_{1}(\mathbf{x}_{0}),\bm{\tilde{t}}_{2}(\mathbf{x}_{0})\right\}$, which approximates a tangent basis of $T_{\mathbf{x}}M$, as well as the last column of $\mathbf{V}_{L}$, denoted by $\bm{\tilde{n}}({\mathbf{x}_{0}})$, which approximates a normal vector (Fig. 1(c)).
@@
    Let $\bar{q}(\tilde{\theta}_{1},\tilde{\theta}_{2})=\sum_{0\leq\alpha_{1}+\alpha_{2}\leq l}\tilde{a}_{\alpha_{1},\alpha_{2}}\tilde{\theta}_{1}^{\alpha_{1}}\tilde{\theta}_{2}^{\alpha_{2}}$ be the optimal solution to the GMLS problem:
@@
In the right panel of Fig. 1, we numerically verify the convergence rate for the estimation of the normal vector $\bm{\hat{n}}$ on a torus where the parametrization of the torus is defined as in (37). The points $\{\mathbf{x}_{i}\}_{i=1}^{N}$ are randomly sampled with a uniform distribution in intrinsic coordinates. One can clearly see that the convergence rate is in good agreement with the above GMLS theoretical error bound of $O(N^{-{l}/{2}})$ in (7) since the normal $\bm{\hat{n}}$ and the two tangent $\bm{\hat{t}}_{1},\bm{\hat{t}}_{2}$ involve with the first-order derivative of the parametrized surface $\bar{q}(\tilde{\theta}_{1},\tilde{\theta}_{2})$.

### v2

@@ Abstract
In this paper, we extend the Generalized Moving Least-Squares (GMLS) method in two different ways to solve the vector-valued PDEs on unknown smooth manifolds without boundaries, identified with randomly sampled point cloud data. The two approaches are referred to as the intrinsic method and the extrinsic method. For the intrinsic method which relies on local approximations of metric tensors, we simplify the formula of Laplacians and covariant derivatives acting on vector fields at the base point by calculating them in a local Monge coordinate system. On the other hand, the extrinsic method formulates tangential derivatives on a submanifold as the projection of the directional derivative in the ambient Euclidean space onto the tangent space of the submanifold. One challenge of this method is that the discretization of vector Laplacians yields a matrix whose size relies on the ambient dimension. To overcome this issue, we reduce the dimension of vector Laplacian matrices by employing a coordinate transformation. The complexity of both methods scales well with the dimension of manifolds rather than the ambient dimension. We also present supporting numerical examples, including eigenvalue problems, linear Poisson equations, and nonlinear Burgers’ equations, to examine the numerical accuracy of proposed methods on various smooth manifolds.
@@ 1 Introduction
Solving vector-valued Partial Differential Equations (PDEs) on curved surfaces or embedded manifolds has various applications in natural sciences and practical engineering. Examples include computation of parallel transport of tangent vectors [53], anisotropic texture synthesis [13, 63], simulation of the magnetohydrodynamics [26], surface hydrodynamics flows [43, 21], surface dynamics of fluid vesicles and biomembranes [4], fluid flows in computer graphics [56].
@@
For aforementioned applications, there have been several developing numerical methods to approximate the solution of vector-valued PDEs on manifolds. Both Discrete Exterior Calculus (DEC) [25, 10] and Finite Element Exterior Calculus (FEEC) [2, 27, 19] are suitable for discretizing the Hodge Laplacian on manifolds. In contrast, the trace FEM [23] was developed to address the Bochner Laplacian and a vector Laplacian associated with the surface strain tensor. Given a sufficient high-quality simplicial complex or mesh, those methods can achieve the desirable convergence in PDE problems for surfaces embedded in $\mathbb{R}^{3}$. However, constructing meshes on surfaces with high quality can be difficult and complicated when the manifold is only identified by randomly sampled data.
@@
To avoid this issue, various mesh-free approaches have been developed for solving PDEs in recent decades. For unknown manifolds with only a point cloud given, the first step of most mesh-free approaches is to characterize the manifolds using a parametrization scheme, such as local SVD methods [12, 66, 62], level set methods [6, 20, 65, 1], closest point methods [49, 45], orthogonal gradient methods [46], moving least squares [37], just to name a few. The second step is to approximate the differential operators acting on functions in the strong form along the tangent bundles using approaches including the Radial Basis Function (RBF) methods [16, 17], Generalized Moving Least-Squares (GMLS) [37, 22], Generalized Finite Difference Methods (GFDM) [59, 57]. In our previous work [24], the RBF method was extended to provide an accurate approximation of differential operators acting on vector fields of smooth manifolds. However, its effectiveness depends crucially on the choice of shape parameters in the radial basis function. Additionally, the computational cost of the method is high since the resulting discrete operators are dense matrices with their size dependent on the ambient dimension. On the other hand, a local mesh method was proposed in [32] which discretizes the weak formulation of differential operators using local connectivity between points. This approach was recently extended to handle vector Laplacians on manifolds in [44].
@@
Alternatively, graph-based approaches, including Vector Diffusion Maps (VDM) [54] and Spectral Exterior Calculus (SEC) [5] can also be used as a mesh-free tool to approximate vector Laplacians for processing raw data in manifold learning applications. While these methods do not require a parametrization of the manifold, it is limited to estimating the Laplacian-type operators.
@@
In this paper, we consider solving vector-valued PDEs on $d$-dimensional unknown manifolds $M$ without boundaries embedded in the Euclidean space $\mathbb{R}^{n}$, identified with randomly sampled point cloud data. Our goal here is to construct an approximation for differential operators acting on vector fields of smooth manifolds where the resulting matrices are sparse with size dependent of the intrinsic dimension $d$ and the number of data $N$. Basically, we translate the formulations from differential geometry into representations that can be concretely and efficiently realized by vector and matrix algebra. Based on two different formulations of vector differential operators, we extend the GMLS methods in two distinct ways to approximate these operators and subsequently solve vector-valued PDEs on manifolds. These two approximations will be referred to as the intrinsic method and the extrinsic method. Our framework is quite flexible for approximating different differential operators such as covariant derivative and various vector Laplacians on simple smooth manifolds.
@@
For the intrinsic GMLS method, it generally involves the local characterization of the Riemannian metric tensor and other geometric quantities of manifolds in local intrinsic coordinates [37, 22, 29]. In this paper, we derive the formulas of the covariant derivative and Laplacians acting on vector fields at the base point in a local Monge coordinate system. It turns out that the resulting formula can be much simplified by using the properties of the Monge coordinate system so that it helps to reduce the computational cost in constructing the vector Laplacian matrices. Basically we find appropriate local and global representations of vector/tensor fields, apply GMLS to each of their associated components, and then assemble these components into a $dN\times dN$ discrete vector Laplacian matrix. Here and throughout, $N$ denotes the size of the point cloud data , $n$ is the ambient dimension and $d$ is the intrinsic dimension of the manifold. Our derivations can be viewed as a generalization from the Laplace-Beltrami of functions in [37] to the covariant derivative and Laplacians of vector fields.
@@
For the extrinsic GMLS method, the surface derivative in the local coordinate can be written as a projection of the appropriate derivative in the ambient space [58, 57, 28]. Then the extrinsic method relies on the discretization in the ambient space which can be computationally expensive for the vector field case. In this paper, we apply a coordinate transformation method to reduce the matrix dimensionality of discrete vector differential operators from $nN\times nN$ to $dN\times dN$. This dimension reduction can be numerically realized since all objects (e.g., functions and tensor fields) in this paper are confined to the tangent bundle of manifolds and the surfaces are all static without movement in the normal direction. In fact, this work improves the result in our previous work [24], saving the memory cost and computational cost especially when the ambient dimension $n\sim 10$ is moderately large.
@@
The paper is organized as follows. In Section 2, we provide a short review of the GMLS approach and its application in the approximation of tangent vectors on unknown manifolds identified by randomly sampled data points. In Section 3, we present the intrinsic formulation for approximating the Bochner Laplacian and the covariant derivative using GMLS. We distinguish our discussion between the cases of 2D surfaces in $\mathbb{R}^{3}$ and general $d$-dimensional manifolds in $\mathbb{R}^{n}$. In Section 4, using the extrinsic formulation, we apply a coordinate transformation to enhance the efficiency of approximation for the Bochner Laplacian and the covariant derivative. In Section 5, we provide supporting numerical examples for solving vector-valued PDEs, particularly eigenvalue problems, screened Poisson equations, linear vector diffusion equations and nonlinear Burgers’ equations. In Section 6, we conclude the paper with a summary and some open problems. We relegate the intrinsic approximation and extrinsic approximation of two other vector Laplacians to the supplementary material Section A and Section B, respectively.
@@ 2.1 Review of GMLS
The GMLS approach is a regression technique for approximating functions by a linear combination of local basis functions, such as local polynomials, based on a set of scattered data samples (see e.g. [42, 38, 35, 64, 39]). We now briefly review the application of GMLS approach on a 2D smooth manifold embedded in $\mathbb{R}^{3}$. Given a set of (distinct) nodes $\mathbf{X}_{M}=\{\mathbf{x}_{i}\}_{i=1}^{N}\subset M$. For a base point ${\mathbf{x}_{0}}\in\mathbf{X}_{M}\subset M$, we denote its $K$-nearest neighbors in $\mathbf{X}_{M}$ by $S_{{\mathbf{x}_{0}}}=\{{\mathbf{x}_{0,k}}\}_{k=1}^{K}\subset\mathbf{X}_{M}$ (Fig. 1(b)). In most literature, such a set $S_{{\mathbf{x}_{0}}}$ is called a “stencil”. By definition, we have ${\mathbf{x}_{0,1}}={\mathbf{x}_{0}}$ to be the base point. We first assume that the true tangent space of the manifold $M$ is given and then discuss the unknown manifold regime in the subsection 2.2. Let $T_{\mathbf{x}}M$ be the tangent space at point $\mathbf{x}=(x^{1},x^{2},x^{3})\in M$ and denote its orthonormal basis by $\left\{\bm{t}_{1}(\mathbf{x}),\bm{t}_{2}(\mathbf{x})\right\}$.
Figure 1: Sketch of our setups. (a) Global tangent-vector bases $\{\bm{t}_{1},\bm{t}_{2}\}$ on the point cloud $\{\mathbf{x}_{i}\}_{i=1}^{N}$ (gray dots) without alignment. (b) $K$-nearest neighbors of the base point $\mathbf{x}_{0}$ plotted as blue crosses inside the red circle. (c) The tangent space $\mathrm{span}\{\bm{t}_{1},\bm{t}_{2}\}$ and the normal $\bm{n}$ at the base point $\mathbf{x}_{0}$. Also plotted are their estimations $\{\bm{\hat{t}}_{1},\bm{\hat{t}}_{2},\bm{\hat{n}}\}$ given in (8) when the manifold is unknown. (d) The local Monge coordinate $(\theta_{1},\theta_{2})$ and the GMLS approximation of the manifold $\hat{q}(\theta_{1},\theta_{2})$ in the local coordinate system (equation (10)).
@@
where $\lambda_{k}$ is some positive weight function. There are various choices of weighted $\ell^{2}$-norms in the moving least-squares (2) as shown in [64, 37, 22]. As reported in previous works [36, 37, 28], the weight function
@@
can numerically provide a stable approximation to the Laplace-Beltrami operator across various data sets [36, 37, 28]. Here we also apply such a special weight function for the numerically stable vector Laplacians. For the covariant derivative, we choose the uniform weight function for a stable approximation.
@@
We now present the theoretical error estimates [39] for such a GMLS approximation for partial derivative operators. For a set of points $\mathbf{X}_{M}=\{\mathbf{x}_{i}\}_{i=1}^{N}\subset M$, define the fill distance $h_{\mathbf{X},M}$ and the separation distance $q_{\mathbf{X},M}$ in Euclidean space as
@@
respectively, and assume that $q_{\mathbf{X},M}\leq h_{\mathbf{X},M}\leq c_{q}q_{\mathbf{X},M}$ for some constant $c_{q}\geq 1$. Let $\mathbf{X}_{M}\subseteq M$ be such that $h_{\mathbf{X},M}\leq h_{0}$ for some constant $h_{0}>0$ and $M^{\ast}=\bigcup_{\mathbf{x}\in M}B(\mathbf{x},C_{2}h_{0})$, a union of geodesic ball over the length $C_{2}h_{0}$ with some constant $C_{2}>0$. Here, the basic idea is to guarantee both the local Taylor expansion and the local polynomial fitting to be valid in the local stencil $B(\mathbf{x},C_{2}h_{0})$ for each $\mathbf{x}\in M$. Then for any $f\in C^{l+1}(M^{\ast})$ and $\bm{\delta}=(\delta_{1},\delta_{2})$ that satisfies $|{\bm{\delta}}|=\delta_{1}+\delta_{2}\leq l$,
@@
for all $\mathbf{x}\in M$ and some $C>0$ (see Section 4 of [39]). In the error bound above, we used the notation $\widehat{D^{{\bm{\delta}}}f}$ as the GMLS approximation to $D^{{\bm{\delta}}}f$ using local polynomials up to degree $l$, where $D^{{\bm{\delta}}}$ denotes a general multi-dimensional derivative with multi-index ${\bm{\delta}}$. For uniformly i.i.d. random data $\mathbf{X}_{M}\subseteq M$ on manifolds, with probability higher than $1-\frac{1}{N}$, one can show that $h_{\mathbf{X},M}=O((\log N/N)^{1/2})$ for a 2D manifold [28]. Consequently, one has the error bound for a two dimensional manifold that $|D^{{\bm{\delta}}}f(\mathbf{x})-\widehat{D^{{\bm{\delta}}}f}(\mathbf{x})|=O\left(\left({\log N}/{N}\right)^{\frac{l+1-|{\bm{\delta}}|}{2}}\right)$.
Remark 2.1.
For a $d$-dimensional manifold, one has $h_{\mathbf{X},M}=O((\log N/N)^{1/d})$ [28]. Utilizing the analysis presented in [39], the following error bound can be established:
$|D^{{\bm{\delta}}}f(\mathbf{x})-\widehat{D^{{\bm{\delta}}}f}(\mathbf{x})|=O\left(\left(\frac{\log N}{N}\right)^{\frac{l+1-|{\bm{\delta}}|}{d}}\right).$ (7)
Although the analysis in [39, 28] was not originally developed for vector fields in the manifold setting, our numerical results (Section 5) show that the convergence rates are in good agreement with the error bound (7).
@@ 2.2 Approximation of tangent vectors for unknown manifolds
For the unknown manifold setup, we assume that we are only given a point cloud $\{\mathbf{x}_{i}\}_{i=1}^{N}\subset M$ on the interior of the smooth manifold. To construct approximations of differential operators, we need to approximate bases of tangent vector space $\left\{\bm{t}_{1}(\mathbf{x}),\bm{t}_{2}(\mathbf{x})\right\}$ at each point $\{\mathbf{x}_{i}\}_{i=1}^{N}$. We first use the local singular value decomposition (SVD) or local principal component analysis (PCA) for a coarse approximation of the tangent vectors [12, 66, 62, 37, 24]:
@@
    Take a singular value decomposition of $\mathbf{D}=\mathbf{V}_{L}\bm{\Sigma}\mathbf{V}_{R}$, and then obtain the leading 2-columns of $\mathbf{V}_{L}\in\mathbb{R}^{3\times 3}$, denoted by $\left\{\bm{\tilde{t}}_{1}(\mathbf{x}_{0}),\bm{\tilde{t}}_{2}(\mathbf{x}_{0})\right\}$, which approximates a tangent basis of $T_{\mathbf{x}_{0}}M$, as well as the last column of $\mathbf{V}_{L}$, denoted by $\bm{\tilde{n}}({\mathbf{x}_{0}})$, which approximates a normal vector (Fig. 1(c)).
@@
    Let $\bar{q}(\tilde{\theta}_{1},\tilde{\theta}_{2})=\sum_{1\leq\alpha_{1}+\alpha_{2}\leq l}\tilde{a}_{\alpha_{1},\alpha_{2}}\tilde{\theta}_{1}^{\alpha_{1}}\tilde{\theta}_{2}^{\alpha_{2}}$ be the optimal solution to the GMLS problem:
@@
    Then $\{\hat{\bm{t}}_{1}(\mathbf{x}_{0}),\hat{\bm{t}}_{2}(\mathbf{x}_{0})\}$ is orthonormalized using Gram-Schmidt process.
@@
Remark 2.2.
For a $d$ dimensional manifold $M$ embedded in $n$ dimensional ambient spaces $\mathbb{R}^{n}$ ($d\ll n$), a similar strategy can be used to approximate the $d$ tangent vectors $\{\bm{\hat{t}}_{i}\}_{i=1}^{d}$ at a given point. Specifically, we first apply the local SVD as described in above Step 1 and Step 2 to obtain a coarse approximation of the tangent vectors $\{\bm{\tilde{t}}_{i}\}_{i=1}^{d}$. Next, we define the local coordinates $\tilde{\theta}_{i}({\mathbf{x}})=\bm{\tilde{t}}_{i}({\mathbf{x}_{0}})\cdot(\mathbf{x}-{\mathbf{x}_{0}}),1\leq i\leq d$. Rather than applying the GMLS approximation to the local normal coordinate $\tilde{q}_{i}(\mathbf{x})=\tilde{\bm{n}}_{i}(\mathbf{x}_{0})\cdot(\mathbf{x}-\mathbf{x}_{0}),1\leq i\leq n-d,$ as done in above 2D case, we instead apply the GMLS to each ambient coordinate of $\mathbf{x}-\mathbf{x}_{0}$. This approach avoided the need for generating $n-d$ normal directions $\tilde{\bm{n}}_{i}(\mathbf{x}_{0})$ and computing $n-d$ inner products in each $\tilde{q}_{i}(\mathbf{x})$ at each point, significantly reducing computational cost when $d\ll n$. A GMLS approximation of $\{\bm{\hat{t}}_{i}\}_{i=1}^{d}$ with higher accuracy can then be calculated as follows:
  1. 3.
    Denote ${\mathbf{x}}=(x^{1},\ldots,x^{n})$ and ${\mathbf{x}}_{0}=(x_{0}^{1},\ldots,x_{0}^{n})$. Let
    $\bar{y}^{s}(\tilde{\theta}_{1},\ldots,\tilde{\theta}_{d})=\bar{x}^{s}(\tilde{\theta}_{1},\ldots,\tilde{\theta}_{d})-x_{0}^{s}=\sum_{1\leq\left|\bm{\alpha}\right|\leq l}\tilde{b}_{\bm{\alpha}}^{s}\bm{\tilde{\theta}}^{\bm{\alpha}}:=\sum_{1\leq\left|\bm{\alpha}\right|\leq l}\tilde{b}_{\alpha_{1},\ldots,\alpha_{d}}^{s}\tilde{\theta}_{1}^{\alpha_{1}}\cdots\tilde{\theta}_{d}^{\alpha_{d}},$
    be the optimal solution of the GMLS problem to approximate $y^{s}=x^{s}-x_{0}^{s}$ for each dimension $s=1,\ldots,n$. Here $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{d})$ and $\left|\bm{\alpha}\right|=\sum_{i=1}^{d}\alpha_{i}$.
  2. 4.
    Solve the normal equation to obtain the regression coefficients $\{\tilde{b}_{\bm{\alpha}}^{s}\}_{1\leq\left|\bm{\alpha}\right|\leq l}$for each $s=1,\ldots,n$. The complexity is $O(nKm^{2}+nm^{3})$ at each ${\mathbf{x}}_{0}$, where $m=\left(\begin{array}[]{c}l+d\\ d\end{array}\right)$ is the number of monomial basis functions $\left\{\bm{\tilde{\theta}}^{\bm{\alpha}}|0\leq\left|\bm{\alpha}\right|\leq l\right\}$.
  3. 5.
    Compute the $d$ tangent vectors of the parametrized manifold $\bar{x}^{s}(\tilde{\theta}_{1},\ldots,\tilde{\theta}_{d})$ at the base $\mathbf{x}_{0}$ in the Cartesian coordinate system as:
    $\bm{\hat{t}}_{k}({\mathbf{x}_{0}})=\sum_{s=1}^{n}\frac{\partial\bar{x}^{s}(0,\ldots,0)}{\partial\tilde{\theta}_{k}}\mathbf{e}_{s}=\sum_{s=1}^{n}\tilde{b}_{\bm{\alpha}(k)}^{s}\mathbf{e}_{s},$ (9)
    where $\bm{\alpha}(k)=(0,\ldots,1,0,\ldots,0)$ with only the $k$th entry being $1$ and $\mathbf{e}_{s}$ is the $s$th standard orthonormal basis of the Cartesian coordinate system. Then we orthonormalize $\{\hat{\bm{t}}_{1}(\mathbf{x}_{0}),\ldots,\hat{\bm{t}}_{d}(\mathbf{x}_{0})\}$ using Gram-Schmidt process.
Notably, we do not need to generate the $n-d$ normal vectors $\{\hat{\bm{n}}_{i}\}_{i=1}^{n-d}$, which would require a computational complexity of about $O(n^{3})$ when $d\ll n$. Now, the total complexity for Step 3 to Step 5 in the above algorithm is $O(n)$. We also point out that the moving least-squares based approach proposed in [55] can be used to handle point cloud data in $\mathbb{R}^{n}$.
(a) 2D Torus in $\mathbb{R}^{3}$ (b) 2D Torus in $\mathbb{R}^{9}$ (c) 3D Flat Torus in $\mathbb{R}^{12}$
Refer to caption Refer to caption Refer to caption
Figure 2: Convergence results for the estimations of projection matrices $\hat{\mathbf{P}}$ on various manifolds. (a) 2D torus in $\mathbb{R}^{3}$. (b) 2D torus in $\mathbb{R}^{9}$. (c) 3D flat torus in $\mathbb{R}^{12}$. Here, $\|\cdot\|_{\mathrm{F}}$ denotes the Frobenius matrix norm, the projection matrix is ${\mathbf{P}}=\sum_{i=1}^{d}\bm{t}_{i}\bm{t}^{\top}_{i}$ and its estimation is $\hat{\mathbf{P}}=\sum_{i=1}^{d}\bm{\hat{t}}_{i}\bm{\hat{t}}^{\top}_{i}$ (see Section 4.1 or [24] for more about the projection matrix).
In Figs. 2(a)-(c), we numerically verify the convergence rate for the estimation of the projection matrices $\hat{\mathbf{P}}=\sum_{i=1}^{d}\bm{\hat{t}}_{i}\bm{\hat{t}}^{\top}_{i}$ on various examples, including a 2D torus in $\mathbb{R}^{3}$, a 2D torus in $\mathbb{R}^{9}$ and a 3D torus in $\mathbb{R}^{12}$. Their parametrizations are defined in (48), (50), and (51), respectively. As shown in Fig. 2, the convergence rate is approximately $O(N^{-{l}/{d}})$, which aligns well with the theoretical GMLS error bound of $O((\log N/N)^{{l}/{d}})$ in (7).